spammmm:

−5−3x=13 i need help with this

MidniteSunny:

alright well first we gotta isolate ze variable (3x) so in order to do dat we'll hav to add 5 from both sides (to cancel out the -5 on the left and balance it out on the right)

Stellasspammm:

Ohhh see thats what i did but then the answer came out wrong so i wasn't sure about what i did wrong

MidniteSunny:

ic, well ze next step ish to divide both sides by 3

Stellasspammm:

ok i did that

MidniteSunny:

alright, wut did u get ?

Stellasspammm:

wait no

Stellasspammm:

6?

MidniteSunny:

o wait sorry i meant divide by -3, mah bad

Stellasspammm:

okok

MidniteSunny:

mhm so wut do u get for that ?

Stelaaaa:

i got 13

MidniteSunny:

hmn nu that's not the correct answer, how did u get that ? did u follow the steps ?

Stelaaaa:

ohhhh waitt sorry i got 18

MidniteSunny:

okay u added 5, great, now u gotta divide the 18 by -3

MidniteSunny:

and divide the -3x by -3 as well

Stelaaaa:

-6?

MidniteSunny:

correct

Stelaaaa:

Yalll I do my work less gooooo

Stelaaaa:

Thank you for helping me

MidniteSunny:

ofc ^^

Ninjashyper456:

do you mean like this Simplifying 5x + -5 = 3x + 13 Reorder the terms: -5 + 5x = 3x + 13 Reorder the terms: -5 + 5x = 13 + 3x Solving -5 + 5x = 13 + 3x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-3x' to each side of the equation. -5 + 5x + -3x = 13 + 3x + -3x Combine like terms: 5x + -3x = 2x -5 + 2x = 13 + 3x + -3x Combine like terms: 3x + -3x = 0 -5 + 2x = 13 + 0 -5 + 2x = 13 Add '5' to each side of the equation. -5 + 5 + 2x = 13 + 5 Combine like terms: -5 + 5 = 0 0 + 2x = 13 + 5 2x = 13 + 5 Combine like terms: 13 + 5 = 18 2x = 18 Divide each side by '2'. x = 9 Simplifying x = 9